OpenStudy (anonymous):

How do you simplify\[\frac{(x-(1+\sqrt{2}))(x-(1-\sqrt{2})) }{(x-3)(x-1)}\]to find the value of x?

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

Hey Oakley~

OpenStudy (bibby):

ew, work. let me try some stuff

OpenStudy (anonymous):

Lol alright

OpenStudy (bibby):

this'd be a lot easier if it were a+b(a-b) form, hmm

OpenStudy (anonymous):

Heh. :p

OpenStudy (bibby):

and combine like terms, let me double check

OpenStudy (anonymous):

wut.

OpenStudy (bibby):

as in \(-2x -2x\sqrt2\)

OpenStudy (anonymous):

brb looking over all the answers sorry I d/c'd

OpenStudy (bibby):

I think I messed something up, let me work it out onpaper

OpenStudy (anonymous):

okay

OpenStudy (bibby):

yeah I flubbed up

OpenStudy (xapproachesinfinity):

you said you want to find the value of x? are you sure because i don't see any equality sign you mean simplify

OpenStudy (anonymous):

Nope, solve for x

OpenStudy (bibby):

I didn't write teh denominator to save space

OpenStudy (anonymous):

Thanks Oakley

OpenStudy (bibby):

the point remains there's nothing to solve for. we simplified the thing

OpenStudy (anonymous):

oh you want the original equation then?

OpenStudy (anonymous):

that's not the original

OpenStudy (anonymous):

this is just a place where I got stuck after following a tutor's instructions, but the tutor isn't here so

OpenStudy (xapproachesinfinity):

here how i would do it: \[\large \frac{[(x-1)+\sqrt{2}][(x-1)-\sqrt{2}]}{(x-3)(x-1)}=\frac{(x-1)^2-2}{(x-3)(x-1)}\]

OpenStudy (xapproachesinfinity):

and carry the simplification

OpenStudy (bibby):

I'm gonna kill myself

OpenStudy (bibby):

rofl, thanks

OpenStudy (xapproachesinfinity):

welcome:) you did good too

OpenStudy (xapproachesinfinity):

in case you are wondering where i did come up with that it is just grouping things

OpenStudy (kittiwitti1):

I think I'm good, thanks.

OpenStudy (xapproachesinfinity):

welcome :)

OpenStudy (kittiwitti1):

...I wonder if anyone noticed? Also - I should know this by now but I don't. I feel dumb. u_u

OpenStudy (xapproachesinfinity):

in your comment above you said this is not the original question so where the rest of the question?

OpenStudy (kittiwitti1):

That wasn't my comment...? :p

OpenStudy (anonymous):

It's in a packet that I will find later

OpenStudy (xapproachesinfinity):

it was the poster's comment!!

OpenStudy (anonymous):

Anyway, brb. Finding the packet

OpenStudy (xapproachesinfinity):

no i just realized, perhaps that was your curiosity when you asked heheh

OpenStudy (xapproachesinfinity):

nvm haha

OpenStudy (anonymous):

Oh I am gonna hate typing this.

OpenStudy (anonymous):

\[\frac{4}{x^{2}-2x-3}\ge\frac{x}{x-3}-\frac{1}{x+1}\text{, solve for }x\]

OpenStudy (xapproachesinfinity):

what do you think?

OpenStudy (anonymous):

What does that mean? e_e

OpenStudy (xapproachesinfinity):

it means what are your thought on how to solve that?

OpenStudy (anonymous):

Uh... dunno

OpenStudy (anonymous):

I mean I did it but I got stuck somewhere along the line

OpenStudy (kittiwitti1):

\[\frac{4}{x^{2}-2x-3}\ge\frac{x(x+1)-(x-3)}{x^{2}-2x-3}\]

OpenStudy (xapproachesinfinity):

the denominator on the left had side factor to (x-3)(x+1) see if that helps

OpenStudy (kittiwitti1):

\[\frac{4}{x^{2}-2x-3}\ge\frac{x^{2}-3x+3}{x^{2}-2x-3}\]

OpenStudy (xapproachesinfinity):

so...

OpenStudy (xapproachesinfinity):

how about you get rid of that same denominator can we do that

OpenStudy (xapproachesinfinity):

but we have to be careful of the sign

OpenStudy (kittiwitti1):

\[\text{Hypothesized solution }0\ge\frac{x^{2}-2x-1}{x^{2}-2x-3}\text{ after shifting the left side over to the right}\]

OpenStudy (anonymous):

so yeah basically went through all those steps and then got stuck

OpenStudy (xapproachesinfinity):

nvm i gotta go but that solution of urs is interesting how did you do that

OpenStudy (triciaal):

do you have the answer? i just have x = 0 or <1

OpenStudy (anonymous):

The tutor gave me the hypothesized solution @xapproachesinfinity

OpenStudy (anonymous):

It does make logical sense but I get to a certain standpoint and I get stuck.

OpenStudy (triciaal):